the rate of change of momentum of a freely falling object is

Grade:12th pass

1 Answers

Arun
6 years ago
Force,
 by definition, is the rate at which momentum changes with time. What is weight, but a force in the vertical direction. 

Momentum = m*v 

Force = Change in momentum = change in m/change in t * v + m*change in v/change in t 

Since mass doesn't usually change, we typically ignore m/t*v, so the equation comes out to be... 
Force = m * v/t 

And change in velocity divided by the change in time is acceleration 

F = m * a